# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ 1 Introduction
The remainder of the paper is organized as follows. In Section 2, we define the notation and prepare some preliminary results. The characterization of first- and second-order stationary points of problem (1.5) is presented in Section 3. We show that second-order stationary points of problem (1.5) correspond to approximate stationary points of problem (1.4) in Section 4. In Section 5, we derive the characterization of the $r$-factorizability of the objective function $f$ in problem (1.4).
@@ 2 Notation and preliminaries
Throughout this paper, we assume that $1\leq r\leq m\leq n$ in problem (1.5). For a matrix $X\in\mathbb{R}^{m\times n}$, we let $\|X\|_{*}$, $\|X\|_{2}$ and $\|X\|_{F}$ denote its nuclear norm, spectral norm and Frobenius norm, respectively. The $i$-th largest singular value of $X$ is denoted by $\sigma_{i}(X)$ for $i=1,\dots,m$. The vector of singular values is denoted by $\sigma(X)=\begin{bmatrix}\sigma_{1}(X)&\cdots&\sigma_{m}(X)\end{bmatrix}^{\top}$. The set of $n\times n$ orthogonal matrices is denoted by $\mathcal{O}^{n}$. For $x\in\mathbb{R}^{s}$, we denote by $\mathrm{Diag}(x)\in\mathbb{R}^{s\times s}$ the diagonal matrix with $(\mathrm{Diag}(x))_{ii}=x_{i}$ for $i=1,\dots,s$. Moreover, we define $\mathrm{diag}:\mathbb{R}^{s\times s}\to\mathbb{R}^{s}$ to be the adjoint operator of $\mathrm{Diag}$. In this paper, to simplify the presentation, we also use $\widetilde{\rm Diag}$ and $\widetilde{\rm diag}$ to denote the possibly non-square versions of $\mathrm{Diag}$ and $\mathrm{diag}$, respectively. Specifically, for $x\in\mathbb{R}^{s}$, $\widetilde{\rm Diag}(x)$ would be a diagonal matrix whose diagonal part is $x$, which is not necessarily square; the dimension of $\widetilde{\rm Diag}(x)$ is omitted when it can be understood from the context.222 For example, if $R\in\mathbb{R}^{m\times m}$ and $P\in\mathbb{R}^{n\times n}$, then writing $R\,\widetilde{\rm Diag}(x)P^{\top}$ would imply that $\widetilde{\rm Diag}(x)\in\mathbb{R}^{m\times n}$. Also, for $X=[X_{1}~{}~{}X_{2}]\in\mathbb{R}^{m\times n}$ with $X_{1}\in\mathbb{R}^{m\times m}$ and $X_{2}\in\mathbb{R}^{m\times(n-m)}$, we define $\widetilde{\rm diag}(X)=\mathrm{diag}(X_{1})\in\mathbb{R}^{m}$. For $X\in\mathbb{R}^{m\times n}$, we define
$\mathcal{O}_{X}=\{(R,P)\in\mathcal{O}^{m}\times\mathcal{O}^{n}:R\,\widetilde{\rm Diag}(\sigma(X))P^{\top}=X\}.$
@@
The set of nonnegative integers is denoted by $\mathbb{N}_{0}$. For a nonnegative integer $r$, we use $[r]$ to denote the set $\{1,\dots,r\}$; in particular, $[0]:=\emptyset$. The permutation group of order $m$ is denoted by $\mathfrak{P}_{m}$, and the set of $m\times m$ permutation matrices is denoted by $\mathcal{P}_{m}$. Finally, for an $x\in\mathbb{R}$, we let $\lfloor x\rfloor$ denote the largest integer upper bounded by $x$.
@@
Let ${\cal R}$ and ${\cal C}$ be the sets consisting of the indices of the rows and columns of $A$ whose sum is less than $1$, respectively. Clearly, ${\cal R}$ and ${\cal C}$ must be simultaneously empty or nonempty. We modify the matrix $A$ gradually in the following manner: at each step, we select $i\in{\cal R}$ and $j\in{\cal C}$, and enlarge $A_{ij}$ until either the row sum of $i$-th row or the column sum of $j$-th column reaches $1$. Then we update ${\cal R}$ and ${\cal C}$ and repeat this process. Since ${\cal R}$ and ${\cal C}$ are always simultaneously empty or nonempty, our algorithm is well defined. Moreover, after each step, the quantity $|{\cal R}|+|{\cal C}|$ is reduced by at least $1$. Since this number is finite, we must end with ${\cal R}={\cal C}=\emptyset$. Then the resulting matrix, denoted by $B$, is doubly stochastic, and it holds by construction that $B_{ij}\geq A_{ij}$ for all $i$ and $j$. ∎
@@
$\mathrm{tr}(R\begin{bmatrix}A&0\end{bmatrix}P\begin{bmatrix}B\\ 0\end{bmatrix})+\mathrm{tr}(R\begin{bmatrix}C&0\end{bmatrix}P\begin{bmatrix}D\\ 0\end{bmatrix})$
@@
Applying Birkhoff theorem (see, e.g., [1, Theorem 1.2.5]), the matrix $Y$ is a convex combination of permutation matrices, namely, $Y=\sum_{i=1}^{s}\lambda_{i}P_{i}$, where $P_{i}\in\mathcal{P}_{m}$, $\lambda_{i}\geq 0$ for each $i=1,\ldots,s$ with $\sum_{i=1}^{s}\lambda_{i}=1$. Therefore, we see that
$(d^{A})^{\top}Yd^{B}+(d^{C})^{\top}Yd^{D}=\lambda_{i}\sum_{i=1}^{s}(d^{A})^{\top}P_{i}d^{B}+(d^{C})^{\top}P_{i}d^{D}\leq\sup_{E\in\mathcal{P}_{m}}(d^{A})^{\top}Ed^{B}+(d^{C})^{\top}E(d^{D}).$
@@ 3 First- and second-order stationary points of $F_{r}$
In this section, we present characterizations of first- and second-order stationary points of $F_{r}$ in problem (1.5), which will be useful for our study of $r$-factorizability in subsequent sections.
@@
where $s:=\mathrm{rank}(V)$ and hence $\sigma_{1}(V),\dots,\sigma_{s}(V)>0$. Denote by $\hat{u}_{i}$ the $i$-th column of $UQ$ for $i\in[m]$. It then follows from (3.1) that the vectors $\hat{u}_{1}/\sigma_{1}(V),\dots,\hat{u}_{s}/\sigma_{s}(V)$ are orthonormal and that $\hat{u}_{i}=0$ for $i=s+1,\dots,m$. There must exist $m-s$ vectors $r_{s+1},\dots,r_{m}$ so that $R=[\hat{u}_{1}/\sigma_{1}(V),\dots,\hat{u}_{s}/\sigma_{s}(V),r_{s+1},\dots,r_{m}]\in\mathcal{O}^{m}$. By the definition of $R$ and the fact that $\sigma(V)=\sigma(U)$, we have
@@
which implies that $(R,Q)\in\mathcal{O}_{U}$ and thus completes the proof. ∎
@@
    Note that the decomposition $-\nabla h(UV^{\top})=R\,\widetilde{\rm Diag}(d)\,P^{\top}$ in 3.2 is not a singular value decomposition in general because it is possible that $d_{s+1}>\lambda=d_{1}=\cdots=d_{s}$. Nevertheless, the vector $d$ contains all the singular values of $-\nabla h(UV^{\top})$, i.e., $d$ is $\sigma(-\nabla h(UV^{\top}))$ up to a permutation of the entries.
@@
    For a stationary point $(U,V)$ of $F_{r}$, 3.2 shows that $\mathrm{rank}(U)=\mathrm{rank}(V)$ and $UV^{\top}=R[\widetilde{\rm Diag}(\sigma(U))]^{2}P^{\top}=R[\widetilde{\rm Diag}(\sigma(V))]^{2}P^{\top}$ for some $R\in\mathcal{O}^{m}$ and $P\in\mathcal{O}^{n}$. Hence, $\sigma_{i}(UV^{\top})=\sigma_{i}^{2}(U)=\sigma_{i}^{2}(V)$ for all $i\in[m]$.
Proof of 3.2.
@@
To prove the “only if” direction, we assume that $(U,V)$ is a stationary point of $F_{r}$ in (1.5), i.e., 3.2 holds. It then follows from [19, Proposition 2] that $U^{\top}U=V^{\top}V$. Fix a singular value decomposition of $V$:
@@
We next establish an equivalent characterization of second-order stationary points of $F_{r}$ in (1.5). We start by introducing a useful way to partition matrices. Let $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ satisfy $\mathrm{rank}(\bar{U})=\mathrm{rank}(\bar{V})$ (which holds in particular when $(\bar{U},\bar{V})$ is a stationary point of $F_{r}$, thanks to 3.2). Denote this common rank by $s=\mathrm{rank}(\bar{U})=\mathrm{rank}(\bar{V})\leq r$. We can then partition any matrices $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$ into the following block form:
where $U_{11},V_{11}\in\mathbb{R}^{s\times s}$, $U_{12},V_{12}\in\mathbb{R}^{s\times(r-s)}$, $U_{21}\in\mathbb{R}^{(m-s)\times s}$, $V_{21}\in\mathbb{R}^{(n-s)\times s}$, $U_{22}\in\mathbb{R}^{(m-s)\times(r-s)}$, $V_{22}\in\mathbb{R}^{(n-s)\times(r-s)}$. Note that when $\bar{U}$ and $\bar{V}$ are of full rank, i.e., $s=r$, the matrices $U_{12}$, $U_{22}$, $V_{12}$ and $V_{22}$ are null.
Proposition 3.4 (Second-order stationarity).
  1. (i)
  2. (ii)
    For any $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$, it holds that [^3] Here, we use the partition 3.6 with respect to $(\bar{U},\bar{V})$; this is well defined because $\sigma(\bar{U})=\sigma(\bar{V})$ holds in Item (i).
    where $\Sigma=\mathrm{Diag}(\sigma_{1}(\bar{U}),\dots,\sigma_{s}(\bar{U}))\in\mathbb{R}^{s\times s}$, and $D=\widetilde{\rm Diag}(d_{s+1},\dots,d_{m})\in\mathbb{R}^{(m-s)\times(n-s)}$ with $d_{i}$ given in Item (i).
Moreover, if $s=\mathrm{rank}(\bar{U})<r$, then Item (i) and Item (ii) imply that $\|\nabla h(\bar{U}\bar{V}^{\top})\|_{2}\leq\lambda$.
Proof.
A pair $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ is a second-order stationary point of $F_{r}$ if and only if it satisfies both the first- and second-order optimality conditions. By 3.2, the first-order optimality condition is equivalent to Item (i).
We now reformulate the second-order optimality condition: $\nabla^{2}F_{r}(\bar{U},\bar{V})[(U,V),(U,V)]\geq 0$ for all $(U,V)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$. Let $R\in\mathcal{O}^{m}$, $P\in\mathcal{O}^{n}$ and $Q\in\mathcal{O}^{r}$ be orthogonal matrices given in Item (i). Since $(U,V)\mapsto(RUQ^{\top},PVQ^{\top})$ is a bijective linear map on $\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$, the condition is equivalent to that $\nabla^{2}F_{r}(\bar{U},\bar{V})[(RUQ^{\top},PVQ^{\top}),(RUQ^{\top},PVQ^{\top})]\geq 0$ for all $(U,V)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$. Denoting $\bar{X}=\bar{U}\bar{V}^{\top}$ and using [19, Equation (3.14)], we have that
$\,\nabla^{2}F_{r}(\bar{U},\bar{V})[(RUQ^{\top},PVQ^{\top}),(RUQ^{\top},PVQ^{\top})]$
$=$  $\,2\langle R^{\top}\nabla h(\bar{X})P,UV^{\top}\rangle+\lambda(\|U\|_{F}^{2}+\|V\|^{2}_{F})+\nabla^{2}h(\bar{X})[\bar{U}QV^{\top}P^{\top}+RUQ^{\top}\bar{V}^{\top}]^{2}.$
The second-order condition is therefore further equivalent to that for all $(U,V)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$,
where (a) follows from the decomposition for $-\nabla h(\bar{X})$ in Item (i) and the definition of $D$ and $\Sigma$, and (b) from the definition of the blocks in 3.6. This shows that Item (i) and Item (ii) together form an equivalent characterization of the second-order stationary points.
We next prove the second claim under the additional assumption of $s=\mathrm{rank}(\bar{U})<r$. Note that this implies that $U_{22}$ and $V_{22}$ are not null. Hence, we can take $U$ and $V$ in 3.7 to be the matrices with the blocks $U_{22}=[e_{1}~{}0]$ and $V_{22}=[e_{1}~{}0]$ and $U_{11}$, $U_{12}$, $U_{21}$, $V_{11}$, $V_{12}$, $V_{21}$ all being zero matrices to deduce that
$-2\sigma_{1}(D)+2\lambda\geq 0.$
The desired conclusion now follows immediately from the above display and the decomposition of $\nabla h(\bar{U}\bar{V}^{\top})$ in Item (i). ∎
4 Approximate stationary points of $f$
Recall that our ultimate goal is to solve problem (1.4) to global optimality. Problem (1.5) is only a surrogate that is computationally more friendly. In practice, it is customary to invoke a first- or second-order optimization algorithm to solve problem (1.5), which often returns a second-order stationary point [17, 30]. A natural question is therefore when the second-order stationary points $(U,V)$ of $F_{r}$ in (1.5) correspond to the global optima of problem (1.4), through the correspondence $(U,V)\mapsto UV^{\top}$. In general, a second-order stationary point of problem 1.5 may not even correspond to a stationary point of problem (1.4). To see this, suppose that $h$ in (1.4) is a strongly convex function, which implies that $f(\cdot)=h(\cdot)+\lambda\|\cdot\|_{*}$ is also strongly convex and hence problem (1.4) has a unique stationary point $X^{*}$. Assume that $\mathrm{rank}(X^{*})>1$ and pick $r<\mathrm{rank}(X^{*})$. Then, $F_{r}$ is the sum of the level-bounded function $(U,V)\mapsto\frac{\lambda}{2}(\|U\|_{F}^{2}+\|V\|_{F}^{2})$ and the function $(U,V)\mapsto h(UV^{\top})$, which is lower bounded due to the strong convexity of $h$. Consequently, the objective function $F_{r}$ is level-bounded, and problem (1.5) must have a global minimizer $(\bar{U},\bar{V})$. However, $\bar{U}\bar{V}^{\top}$ is not a stationary point of $f$, since $\mathrm{rank}(\bar{U}\bar{V}^{\top})\leq r<\mathrm{rank}(X^{*})$. Although in general second-order stationary points of $F_{r}$ do not correspond to a stationary point of $f$ in 1.4, we show that they correspond to an approximate stationary point of $f$. A similar result was obtained in [15, Theorem 1] under the restricted isometry property.
Theorem 4.1 (Approximate stationary points).
Let $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ be a second-order stationary point of $F_{r}$ in 1.5, $s=\mathrm{rank}(\bar{U})$ and $d\in\mathbb{R}^{m}$ be given in Proposition 3.4. Then, [^4] We define $d_{m+1}=0$. $d_{s+1}\leq\lambda+L\sigma_{r}(\bar{U}\bar{V}^{\top})$, where
$L=\sup_{\begin{subarray}{c}Y\in\mathbb{R}^{m\times n},\,\|Y\|_{F}=1,\\ \mathrm{rank}(Y)=2\end{subarray}}\nabla^{2}h(\bar{U}\bar{V}^{\top})[Y,Y].$
Furthermore, $\inf_{Y\in\partial f(\bar{U}\bar{V}^{\top})}\|Y\|_{2}\leq L\,\sigma_{r}(\bar{U}\bar{V}^{\top})$.
Proof.
$d_{s+1}\leq\|d\|_{\infty}=\|\nabla h(\bar{U}\bar{V}^{\top})\|_{2}\leq\lambda=\lambda+L\sigma_{r}(\bar{U}\bar{V}^{\top}),$
where the first equality follows from Remark 3.3(i) and the second inequality follows from 3.4. Therefore, we have
$-\frac{1}{\lambda}\nabla h(\bar{U}\bar{V}^{\top})=R\widetilde{\rm Diag}(1,\dots,1,d_{s+1}/\lambda,\dots,d_{m}/\lambda)P^{\top}\in\partial\|\bar{U}\bar{V}^{\top}\|_{*},$
where the inclusion follows from 2.1 and the fact that $0\leq d_{m}\leq\cdots\leq d_{s+1}\leq\lambda$. This proves $0\in\partial f(\bar{U}\bar{V}^{\top})$, and hence $00=\inf_{Y\in\partial f(\bar{U}\bar{V}^{\top})}\|Y\|_{2}\leq L\,\sigma_{r}(\bar{U}\bar{V}^{\top})=$ 0.
We next consider the case where $\mathrm{rank}(\bar{U})=r$. If $r=m$, then $d_{m+1}=0$; moreover, we have $d_{1}=\cdots=d_{m}=\lambda$ from 3.4, which implies (as in the above display) that $\inf_{Y\in\partial f(\bar{U}\bar{V}^{\top})}\|Y\|_{2}=0$. The desired conclusions then hold trivially. Thus, from now on, we assume $r<m$.
By 3.2, $\mathrm{rank}(\bar{V})=\mathrm{rank}(\bar{U})=r$. Therefore, in this case, the blocks $U_{12},U_{22},V_{12},V_{22}$ in (3.6) are null. Since $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$ in 1.5, by 3.4, it satisfies the following inequality for any matrices $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$:

### v2

@@ 1.1 Sketch of proof techniques
The remainder of the paper is organized as follows. In Section 2, we define the notation and present some preliminary results. The characterization of first- and second-order stationary points of problem (1.2) is presented in Section 3. In Section 4, we prove 1.3 by using the results in Section 2 and Section 3. The generalization to the $\mathrm{RIP}$ case, namely Corollary 1.4, is proved in Section 5.
@@ 2 Notation and preliminaries
Throughout this paper, we assume that $1\leq r\leq m\leq n$ in problem (1.2). For a matrix $X\in\mathbb{R}^{m\times n}$, we let $\|X\|_{*}$, $\|X\|_{2}$ and $\|X\|_{F}$ denote its nuclear norm, spectral norm and Frobenius norm, respectively. The $i$-th largest singular value of $X$ is denoted by $\sigma_{i}(X)$ for $i=1,\dots,m$. The vector of singular values is denoted by $\sigma(X)=\begin{bmatrix}\sigma_{1}(X)&\cdots&\sigma_{m}(X)\end{bmatrix}^{\top}$. The set of $n\times n$ orthogonal matrices is denoted by $\mathcal{O}^{n}$. For $x\in\mathbb{R}^{s}$, we denote by $\mathrm{Diag}(x)\in\mathbb{R}^{s\times s}$ the diagonal matrix with $(\mathrm{Diag}(x))_{ii}=x_{i}$ for $i=1,\dots,s$. Moreover, we define $\mathrm{diag}:\mathbb{R}^{s\times s}\to\mathbb{R}^{s}$ to be the adjoint operator of $\mathrm{Diag}$. In this paper, to simplify the presentation, we also use $\widetilde{\rm Diag}$ and $\widetilde{\rm diag}$ to denote the possibly non-square versions of $\mathrm{Diag}$ and $\mathrm{diag}$, respectively. Specifically, for $x\in\mathbb{R}^{s}$, $\widetilde{\rm Diag}(x)$ would be a diagonal matrix whose diagonal part is $x$, which is not necessarily square; the dimension of $\widetilde{\rm Diag}(x)$ is omitted when it can be understood from the context. [^3] For example, if $R\in\mathbb{R}^{m\times m}$ and $P\in\mathbb{R}^{n\times n}$, then writing $R\,\widetilde{\rm Diag}(x)P^{\top}$ would imply that $\widetilde{\rm Diag}(x)\in\mathbb{R}^{m\times n}$. Also, for $X=[X_{1}~~X_{2}]\in\mathbb{R}^{m\times n}$ with $X_{1}\in\mathbb{R}^{m\times m}$ and $X_{2}\in\mathbb{R}^{m\times(n-m)}$, we define $\widetilde{\rm diag}(X):=\mathrm{diag}(X_{1})\in\mathbb{R}^{m}$. For $X\in\mathbb{R}^{m\times n}$, we define
$\mathcal{O}_{X}:=\{(R,P)\in\mathcal{O}^{m}\times\mathcal{O}^{n}:R\,\widetilde{\rm Diag}(\sigma(X))P^{\top}=X\}.$
@@
The set of nonnegative integers is denoted by $\mathbb{N}_{0}$. For a nonnegative integer $r$, we use $[r]$ to denote the set $\{1,\dots,r\}$; in particular, $0]:=\emptyset$. The permutation group of order $m$ is denoted by $\mathfrak{P}_{m}$, and the set of $m\times m$ permutation matrices is denoted by $\mathcal{P}_{m}$. Finally, for an $x\in\mathbb{R}$, we let $\lfloor x\rfloor$ denote the largest integer upper bounded by $x$.
@@
[…]
@@
[…]
@@
Applying Birkhoff theorem (see, e.g., [2, Theorem 1.2.5]), the matrix $Y$ is a convex combination of permutation matrices, namely, $Y=\sum_{i=1}^{s}\lambda_{i}P_{i}$, where $P_{i}\in\mathcal{P}_{m}$, $\lambda_{i}\geq 0$ for each $i=1,\ldots,s$ with $\sum_{i=1}^{s}\lambda_{i}=1$. Therefore, we see that
$(d^{A})^{\top}Yd^{B}+(d^{C})^{\top}Yd^{D}=\sum_{i=1}^{s}\lambda_{i}[(d^{A})^{\top}P_{i}d^{B}+(d^{C})^{\top}P_{i}d^{D}]\leq\sup_{E\in\mathcal{P}_{m}}(d^{A})^{\top}Ed^{B}+(d^{C})^{\top}E(d^{D}).$
@@ 3 First- and second-order stationary points of $F_{r}$
In this section, we present characterizations of first- and second-order stationary points of $F_{r}$ in problem (1.2). Let us note that in the literature, the stationarity of $F_{r}$ has already been studied in [19, 20]. However, for our purpose, we need to extract useful features that have not been documented in the literature. To begin with, we can derive directly from the first-order optimality condition of problem (1.2) (see (3.2) below; see also [19, Proposition 2]) that if $(U,V)$ is a stationary point of $F_{r}$, then $U^{\top}U=V^{\top}V$, which is also called a balanced pair. The next lemma studies how the balancedness can affect the singular vectors of $U$ and $V$, which serves as the basis to establish that $X$ and $\nabla h(X)$ have SVDs with common orthonormal bases (possibly in different order) when $X=UV^{\top}$ with $(U,V)$ being a first-order stationary point of $F_{r}$.
@@
where $s:=\mathrm{rank}(V)$ and hence $\sigma_{1}(V),\dots,\sigma_{s}(V)>0$. Denote by $\hat{u}_{i}$ the $i$-th column of $UQ$ for $i\in[m]$. It then follows from (3.1) that the vectors $\hat{u}_{1}/\sigma_{1}(V),\dots,\hat{u}_{s}/\sigma_{s}(V)$ are orthonormal and that $\hat{u}_{i}=0$ for $i=s+1,\dots,m$. There must exist msm-s vectors $r_{s+1},\dots,r_{m}$ so that $R=[\hat{u}_{1}/\sigma_{1}(V),\dots,\hat{u}_{s}/\sigma_{s}(V),r_{s+1},\dots,r_{m}]\in\mathcal{O}^{m}$. By the definition of $R$ and the fact that $\sigma(V)=\sigma(U)$, we have
@@
which implies $(R,Q)\in\mathcal{O}_{U}$ and thus completes the proof. ∎
Next, we are ready to establish the SVD structures of $X$ and $\nabla h(X)$ when $X=UV^{\top}$ with $(U,V)$ being a first-order stationary point of $F_{r}$. The proof is basically done by substituting the SVDs of $(U,V)$ into the first-order optimality condition and then solving the stationary equations.
@@
    Note that the decomposition $\nabla h(UV^{\top})=R\,\widetilde{\rm Diag}(d)\,P^{\top}$ in Proposition 3.2 is not a singular value decomposition in general because it is possible that $d_{s+1}>\lambda=d_{1}=\cdots=d_{s}$. Nevertheless, the vector $d$ contains all the singular values of $\nabla h(UV^{\top})$, i.e., $d$ is $\sigma(-\nabla h(UV^{\top}))$ up to a permutation of the entries.
@@
    For a stationary point $(U,V)$ of $F_{r}$, Proposition 3.2 shows that $\mathrm{rank}(U)=\mathrm{rank}(V)$ and $UV^{\top}=R[\widetilde{\rm Diag}(\sigma(U))]^{2}P^{\top}=R[\widetilde{\rm Diag}(\sigma(V))]^{2}P^{\top}$ for some $R\in\mathcal{O}^{m}$ and $P\in\mathcal{O}^{n}$. Hence, $\sigma_{i}(UV^{\top})=\sigma_{i}^{2}(U)=\sigma_{i}^{2}(V)$ for all $i\in[m]$.
Proof of Proposition 3.2.
@@
To prove the “only if” direction, we assume that $(U,V)$ is a stationary point of $F_{r}$ in (1.2), i.e., 3.2 holds. A direct computation shows that $U^{\top}U=V^{\top}V$; see also [19, Proposition 2]. Fix a singular value decomposition of $V$:
@@
In view of the first-order optimality condition of 1.1 and Proposition 2.1, one can see that the SVD structures of $UV^{\top}$ and $\nabla h(UV^{\top})$ given by Proposition 3.2 share some features of the true stationary point $X^{*}$ of $f$ in 1.1 and the corresponding $\nabla h(X^{*})$, even though $UV^{\top}$ is not necessarily a stationary point of $f$. To treat such points and the true stationary points under a unified framework, we define the following category of points of the function $f$.
Definition 3.4 (Pseudo-stationarity).
A matrix $X\in\mathbb{R}^{m\times n}$ is said to be a pseudo-stationary point of $f$ in (1.1) if there exist $(R,P)\in\mathcal{O}_{X}$ and $d\in\mathbb{R}^{m}_{+}$ such that $\nabla h(X)=R\widetilde{\rm Diag}(d)P^{\top}$ and $d_{1}=\cdots=d_{s}=\lambda$, where $s=\mathrm{rank}(X)$.
Assume that $X^{*}$ and $\bar{X}$ are two pseudo-stationary points of $h$. When $h$ is $\mu$-strongly convex and $\nabla h$ is $L$-Lipschitz continuous, we know that 1.8 holds with $X=X^{*}$ and $Y=\bar{X}$. We next leverage Proposition 3.2, which states that $X^{*}$ and $\nabla h(X^{*})$, as well as $\bar{X}$ and $\nabla h(\bar{X})$, admit simultaneous SVDs, to transform 1.8 to a bound involving only the singular values of $X^{*}$, $\nabla h(X^{*})$, $\bar{X}$ and $\nabla h(\bar{X})$.
Lemma 3.5.
Let $X_{1},X_{2}$ be two pseudo-stationary points of $f$ in 1.1 in the sense of Definition 3.4, i.e., there exist $R_{1}$, $R_{2}\in\mathcal{O}^{m}$ and $P_{1}$, $P_{2}\in\mathcal{O}^{n}$ such that
$X_{1}=R_{1}\begin{bmatrix}\Sigma_{1}&0_{m\times(n-m)}\end{bmatrix}P_{1}^{\top},~-\nabla h(X_{1})=R_{1}\begin{bmatrix}D_{1}&0_{m\times(n-m)}\end{bmatrix}P_{1}^{\top},$ (3.6)
where $\Sigma_{i}=\mathrm{diag}(\sigma_{1}(X_{i}),\dots,\sigma_{m}(X_{i}))\in\mathbb{R}^{m\times m}$ and $D_{i}=\mathrm{diag}(d_{1}^{i},\dots,d_{m}^{i})\in\mathbb{R}^{m\times m}_{+}$ with $d^{i}_{1}=\dots=d^{i}_{\mathrm{rank}(X_{i})}=\lambda$ for $i=1,2$. Assume that $h$ in 1.1 satisfies that $h$ is $\mu$-strongly convex for some $\mu>0$, and $\nabla h$ is Lipschitz continuous with modulus $L\geq\mu$. Then, we have